Tim Hortons paid its line workers $10 per hour lest yeer when the Consumer Price Index wes 100 Suppose over the past year, deflation occurred and the aggregate price level fe Tim Hortons paid its line workers $10 per hour last year when the to 80 instructions: Round your answers to two decimal places this year in order to keep the real wage fixed ot $10 a. Tim Hortons must pay its workers s b. Tim Hortons must pay its workers c. If Tim Hortons keeps the wage fixed at $10 per hour, in real terms, its workers get a this year if it wants to increase the real wage by 10 peicent % increase in real wages.
Answer #1

a) Tim paid $10 per hour last year when consumer index was 100 but now because of deflation, it has been down to 80.
In such a case, to keep real wage at that level, he should pay $8 per hour.

( 80 * 10 ) / 100 = 8

b) Now if Tim want to increase real wage by 10% then he must pay $8.8 per hour.
Because of deflation real wage is $8 now.
8 + ( 8 * 10 / 100 ) = 8.8

c) If Tom has fixed the wage at $10 then it is higher by $2 than the real wage of $8.

( 100 * 2 ) / 8 = 25

At $10, there is 25% increase in real wages.